OpenStudy (anonymous):

a particle is projected with velocity 9m/s up a line of greatest slope of a smooth plane inclined at an angle a to the horizontal where sin a=4/7. find the speed of the particle when it has travelled 5m up the plane.

OpenStudy (aravindg):

u get a by taking arcsin 4/7

OpenStudy (anonymous):

then?

OpenStudy (aravindg):

srry i gota do my problems bye i am sure smone else will help

OpenStudy (aravindg):

wel i'll giv u hint

OpenStudy (aravindg):

use s=ut+1/2 at^2 for horizontal component and get time substitute this time in any other eqn to get ur v

OpenStudy (anonymous):

i get the answer 11.71m/s using formula s=vt-1/2at^2. is it correct?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

use equation v^2=u^2+2as

OpenStudy (anonymous):

what is a?

OpenStudy (anonymous):

a=-g(sin a)

OpenStudy (anonymous):

\[-mg \sin \Theta =ma\] sin theta=4.7

OpenStudy (anonymous):

so our formula is matched but final a isn't

OpenStudy (anonymous):

how could??

OpenStudy (anonymous):

a=-g(sin a) & \[v=\sqrt{(v _{0}^{2}+2as)}\]

OpenStudy (anonymous):

then,u get same answer as me or not?

OpenStudy (anonymous):

i distracted no .never mind i'll find it

OpenStudy (anonymous):

unbelievable..

OpenStudy (anonymous):

i think 4.99 is correct hence particles velocity ought be decreased

OpenStudy (anonymous):

okey. thanx a lot.

OpenStudy (anonymous):

ur wellcome friend

OpenStudy (aravindg):

no medals??